Table 1: Pepita Disco Performance, 2011 (UYU in millions)

Units sold 100 million units
Revenue 200
Variable costs
Materials 30
Direct labor(manufacturing,sales) 40
Opearational costs(manufacturing,inventory,delivery) 30
Other variable costs 20
Margin on Sales 80
Fixed Costs
Marketing and advertising 10
Research and development 10
Other fixed costs(e.g head office) 40
Net margin 20

Now, imagine a similar company, called Yuckles Pet Products, which sells more expensive products than Pepita Disco. It has exactly the same revenues, total fixed costs, and total variable costs as Pepita Disco, so its financials look exactly the same except that Yuckles sold 40 million units.

1. What would be the absolute and percentage effect on net margin if Yuckles were to:

a.) Lower its price 10% with an elasticity of –1.7?
Answer #1

1. Current price/product for Yuckles = Total Revenue/Total Units = 200/40 = $ 5/Product

If Yuckles lower its price 10% with an elasticity of –1.7

Revised Price = 5 - (10% of 5) = $ 4.5/product

Demand would increase by = 40 Million * 17% = 46.8 Million

Calculation of variable cost per unit
Units sold (Yuckles) 40 million units Per Unit
Revenue 200 5
Variable costs
Materials 30 0.75
Direct labor(manufacturing,sales) 40 1
Opearational costs(manufacturing,inventory,delivery) 30 0.75
Other variable costs 20 0.5
Revised Profit table for 46.8 million units
Units sold (Yuckles) 46.8 million units Per Unit
Revenue 210.6 4.5
Variable costs
Materials 35.1 0.75
Direct labor(manufacturing,sales) 46.8 1
Opearational costs(manufacturing,inventory,delivery) 35.1 0.75
Other variable costs 23.4 0.5
Margin on Sales 70.2
Fixed Costs
Marketing and advertising 10 NA
Research and development 10 NA
Other fixed costs(e.g head office) 40 NA
Net margin 10.2

Absolute increase in net margin = 10.2 -20 = $ 9.8 Million decrease in margin

Percentage effect = -9.8/20 = 49%